OpenStudy (anonymous):

What is a quartic function with only the two real zeros given? x = -4 and x = -1 A. y = x4 + 5x3 + 5x2 + 5x + 4 B. y = x4 - 5x3 - 5x2 - 5x - 4 C. y = -x4 + 5x3 + 5x2 + 5x + 4 D. y = x4 + 5x3 + 5x2 + 5x - 5

OpenStudy (anonymous):

It seems that \[ x^4+5 x^3+5 x^2+5 x+4=(x+1) (x+4) \left(x^2+1\right) \] What is your choice?

OpenStudy (anonymous):

Thank you...Ummmm they look like this A. y = x4 + 5x3 + 5x2 + 5x + 4 B. y = x4 - 5x3 - 5x2 - 5x - 4 C. y = -x4 + 5x3 + 5x2 + 5x + 4 D. y = x4 + 5x3 + 5x2 + 5x - 5

OpenStudy (anonymous):

Are you factoring?

OpenStudy (anonymous):

Ohhhh is it A

OpenStudy (anonymous):

Yes